a car travels in 270 miles in 4.5 hour. what is the average speed? how far it will go in 7 hours at this average speed?
Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

279/4.5=60 mph. is the average speed.
7*69=420 miles. is the distance driven @ 60 mph. for 7 hours.

Answer by richwmiller(17219)

rt=d
r(4.5)=270 miles
r=270/4.5=60 mph
60*7=420 miles in 7 hours at 60 mph